OpenStudy (anonymous):

Quadratic equations, which are expressed in the form of ax^2+bx+c=0, where a does not equal 0, may have how many solutions…explain

OpenStudy (campbell_st):

well if a is not equal to zero means that the quadratic will have up to 2 solutions... here are some examples \[x^2 - 6x + 9 = 0.... is....(x - 3)^2 = 0 \] has a repeated solution x = 3 or a solution with multiplicity 2 \[x^2 + x - 6 = 0.... is .... (x -2)(x+3) = 0\] has 2 unequal solutions x = 2 and x = -3 \[x^2 + x + 2 = 0\] has no real number solutions.... as the quadratic has complex number solutions.